# Vibe Coding XR: Accelerating AI + XR Prototyping with XR Blocks and Gemini

### 25 Mar 2026 v1

@@ 4. Preliminary Technical Evaluation
Evaluating XR applications poses unique challenges, as it typically requires manual, on-device testing and subjective human evaluation. However, this manual process struggles to keep pace with the rapid iteration cycles of foundational models. To assess the effectiveness of our pipeline at scale, we constructed a preliminary benchmark dataset, VCXR60, to automate testing.
@@
Sourced from four one-hour workshops, VCXR60 consists of 60 XR application prompts provided by 20 participants (detailed in Appendix A). Using this dataset together with the XR Blocks desktop simulator, we measured inference time and the one-shot pass rate (pass@1). Inspired by code LLM benchmarks like HumanEval (Chen et al., 2021), we define a “pass” as zero-error execution via unit test. In 3D graphics, the rendering engine is an inherent unit test engine to leverage, and thus we monitor runtime error logs in a headless Chromium browser via Playwright (Playwright, 2026) as our unit test.
@@
Early analysis revealed that the majority of initial generation errors stemmed from edge-case bugs within the XR Blocks framework itself (before v0.11.0) or hallucinated APIs by the LLMs, initially yielding an approximate 70% pass rate. These insights fueled a rapid, six-month iteration cycle. After 11 major releases, we evaluated XR Blocks v0.11.0 to gather the baseline results shown in Table 1.
@@
Averaged across five runs, our evaluation revealed distinct performance trade-offs based on model size, “thinking” levels, and prompt complexity. Notably, every test prompt achieved at least one successful execution across the five runs. For simpler interactions (e.g., “Create a beautiful dandelion that blows away when I pick it up”), faster models like Gemini Flash often completed generation in under 20 seconds. However, for applications requiring complex animation states and precise hand-interaction logic (e.g., Prompt 060 in Table 1), these models exhibited a higher rate of runtime errors compared to more advanced models like Gemini Pro.
Our primary takeaway for researchers and practitioners is clear: larger models are essential for authoring XR applications that intricately weave perception with interaction. Enabling a “high” thinking mode yields the most reliable results for advanced XR prototyping. By affording the LLM more tokens to explicitly plan its thought process, this workflow more successfully navigates the complex spatial logic, physics integrations, and perception pipelines required for XR, drastically improving the one-shot success rate and minimizing debugging friction for the user.
@@
5. Limitation and Discussion
This paper present our initial step towards “idea to reality”. Our current implementation prioritizes the core abstractions for XR, and its limitations highlight several exciting avenues for future research:
Performance and Latency: Our choice of web technologies prioritizes accessibility but brings known trade-offs. The framework can never match the rendering performance of native engines like Unity or Unreal, and reliance on cloud AI models introduces network latency. Our visionary goal is to develop an LLM-driven cross-compiler capable of translating a high-level XR Blocks script into optimized, native code across multiple target engines like GoDot; we invite the whole community to contribute the SDK for common practice of AI + XR interactions to further reduce LLM token costs.
Multi-Sensory Synthesis: We have focused on audio-visual experiences, but XR Blocks is designed to be extensible. Integrating modules for haptics, EEG, and other sensory modalities could allow creators to compose truly immersive, multi-sensory narratives. E.g., one may extend XR Blocks with Arduino through WebUSB protocol.
Large-scale Benchmarking and Agentic Evaluation: As an initial step, VCXR-60 is far from covering the full spectrum of AI + XR. For instance, a rigorous evaluation should decouple contributions of XR Blocks from three.js, and quantitative evaluate how many tokens XR Blocks could save with the abstradctions. While we are committed to developing a more rigorous benchmark, we would like to evolve XR Blocks together with the next generation of LLMs by open-sourcing with the community. Future work will look into agentic testing with human-in-the-loop and eventually lead to inventing better interaction paradigms for AI + XR.
@@ 6. Conclusion
Ultimately, we open-source this framework as an open invitation. We welcome the HCI, XR, and AI communities to build upon this workflow, fill the existing gaps in tooling primitives and expressivity, and collectively realize a future where creating an interactive, mixed-reality world is as natural as describing a dream.
@@ Acknowledgements.
We thank all of XR Blocks contributers in 2025, including David Li, Nels Numan, Xun Qian, Yanhe Chen, Zhongyi Zhou, Evgenii Alekseev, Geonsun Lee, Alex Cooper, Min Xia, Scott Chung, Jeremy Nelson, Xiuxiu Yuan, Jolica Dias, Tim Bettridge, Benjamin Hersh, Michelle Huynh, Konrad Piascik, Ricardo Cabello, David Kim, and Ruofei Du. We further thank Xiuxiu Yuan, Oren Haskins, Brian Collins, Barak Moshe, Seeyam Qiu, and Coco Fatus for the UX Design Support; Yinghua Yang and Brenton Simpson for the cross-team collaboration; Tim Bettridge, Kevin Lam for significant contributions to XR Blocks x Gemini Canvas; Furthermore, we would like to extend thanks to the seminal work in Android XR Unity Samples with Mahdi Tayarani, Max Dzitsiuk, Patrick Hackett, all of DepthLab (Du et al., 2020), Ad hoc UI (Du et al., 2022), Rapsai (Visual Blocks) (Du et al., 2023), InstructPipe (Zhou et al., 2025), DialogueLab (Hu et al., 2025a), and Sensible Agent (Lee et al., 2025) co-authors, which greatly inspired this project along the way; we further extend thanks to Jim Ratcliffe, Eric Gonzalez, Nicolás Peña Moreno, Max Spear, Yi-Fei Li, Ziyi Liu, Jing Jin, Sean Fanello, Thabo Beeler, Adarsh Kowdle, and Guru Somadder for their insightful feedback and support in XR Blocks.References

### 03 Apr 2026 v2

@@ 4. Preliminary Evaluation with VCXR60
Evaluating XR applications poses unique challenges, as it typically requires manual, on-device testing and subjective human evaluation. However, assessing the viability of generative XR systems requires fundamentally new, scalable technical metrics that can keep pace with the rapid iteration of large language models. To meet this need, we introduce the VCXR60 dataset, an open-source benchmark designed to automate the testing of intent-driven spatial computing.
@@
Sourced from four one-hour workshops, VCXR60 consists of 60 XR application prompts provided by 20 participants (detailed in Appendix A). Using this dataset with XR Blocks’s desktop simulator, we measured inference time and one-shot pass rates (pass@1). Inspired by code LLM benchmarks like HumanEval (Chen et al., 2021), we define a “pass” as a zero-error execution in an automated runtime test. In 3D graphics, the rendering engine itself serves as an implicit correctness validator, as runtime errors during rendering indicate code failures. We therefore monitor error logs in a headless Chromium browser via Playwright (Playwright, 2026) as our automated test harness.
@@
Early analysis revealed that the majority of generation errors stemmed from edge-case bugs within the XR Blocks framework itself (v0.1.0) or hallucinated APIs by the LLMs, initially yielding an approximate 70% pass rate. These insights informed a rapid six-month iteration cycle. After 10 major releases, we evaluated XR Blocks v0.11.0 to gather the baseline results shown in Table 1.
@@
Averaged across five runs, our evaluation revealed distinct performance trade-offs based on model size, “thinking” levels, and prompt complexity. Notably, every test prompt achieved at least one successful execution across the five runs. For simpler interactions (e.g., “Create a beautiful dandelion that blows away when I pick it up”), faster models like Gemini Flash often completed generation in under 20 seconds. For applications requiring complex animations and precise hand-interaction logic (e.g., Prompt 060 in Appendix A), however, these models exhibited a higher rate of runtime errors compared to more advanced models like Gemini Pro.
@@
5. Limitations and Discussion
This paper presents our initial step towards “idea to reality”. Our current implementation prioritizes the core abstractions for XR, and its limitations highlight exciting avenues for future research.
Reliability of LLM-Generated Code
The stochastic nature of LLMs introduces friction into the prototyping loop. Generated applications occasionally misinterpret physical constraints, produce invalid syntax, or fail to implement complex interaction logic. Our evaluation shows that larger models with extended reasoning budgets mitigate these issues considerably, but do not eliminate them. A systematic investigation of failure modes, particularly for prompts that require tightly coupled perception and interaction logic, would help identify where additional framework-level abstractions or constrained generation strategies could improve reliability.
Performance and Platform Trade-offs
Our choice of web technologies prioritizes accessibility and shareability but introduces known trade-offs. The framework cannot match the rendering throughput of native engines like Unity or Unreal, and reliance on cloud-hosted LLMs adds network latency to the generation step. A longer-term goal is to develop an LLM-driven cross-compiler capable of translating high-level XR Blocks scripts into optimized, native code for target engines such as Godot (Godot, 2022). We invite the community to contribute to the SDK to establish shared conventions for AI + XR interactions and to further reduce LLM token costs.
Toward Rigorous Benchmarking
VCXR60 covers a limited portion of the AI + XR design space, and our automated pass/fail metric captures code viability but does not assess usability, aesthetic quality, or interaction fidelity. A rigorous evaluation should decouple the contributions of XR Blocks from those of the underlying three.js library and quantify how many tokens the framework’s abstractions save relative to raw API usage. Evaluation should also address iterative, multi-turn usage scenarios, such as refining an existing prototype in response to evolving requirements, which may expose distinct failure modes in context management and incremental code coherence. Just as ImageNet (Deng et al., 2009) accelerated progress in computer vision, the XR and HCI fields would benefit from large-scale, open-source datasets of prompt-to-interaction pairs that enable systematic evaluation of generated spatial behaviors through human raters, expert review, and agentic testing workflows. We aim to evolve XR Blocks alongside advances in LLMs through continued open-source collaboration with communities. For example, XR Blocks currently lacks ready-to-use accessibility and safety modules, an important gap given the framework’s goal of broad participation. Future iterations of Vibe Coding XR would benefit from rigorous human-in-the-loop evaluation to understand how designers co-create with AI agents.
@@ 6. Conclusion
Ultimately, we release this framework, workflow, and preliminary dataset as an open invitation. We welcome the HCI, XR, and AI communities to build upon this foundation, expand the capabilities of generative spatial computing, and collectively work toward a future where moving from “idea to reality” is as fluid as describing it, and eventually kicking off the flywheel of accelerating AI + XR innovations together.
@@ Acknowledgements.
We thank all of XR Blocks contributers in 2025: David Li and Ruofei Du (equal primary contributions), Nels Numan, Xun Qian, Yanhe Chen, and Zhongyi Zhou, (equal secondary contributions, sorted alphabetically), as well as Evgenii Alekseev, Geonsun Lee, Alex Cooper, Brandon Jones, Min Xia, Scott Chung, Jeremy Nelson, Xiuxiu Yuan, Jolica Dias, Tim Bettridge, Benjamin Hersh, Michelle Huynh, Konrad Piascik, Ricardo Cabello, and David Kim. We further thank the Gemini Canvas and AI Studio teams for their support including, but not limited to: Tim Bettridge, Yan Li, Daniel Marques, Deven Tokuno, Levent Yilmaz, Saravana Rathinam, Samuel Petit, Mike Taylor-Cai, Ammaar Reshi, and Robert Berry, We would like to thank Mahdi Tayarani, Max Dzitsiuk, Jim Ratcliffe, Patrick Hackett, Seeyam Qiu, Coco Fatus, Alon Hetzroni, Aaron Kim, Yinghua Yang, Brian Collins, Eric Gonzalez, Keith Moon, Nicolás Peña Moreno, Yidang Zhang, Jamie Pepper, Yuhao He, Yi-Fei Li, Ziyi Liu, Jing Jin for their feedback and discussion on our early-stage proposal and WebXR experiments, as well as all of DepthLab (Du et al., 2020), Ad hoc UI (Du et al., 2022), Rapsai (Visual Blocks) (Du et al., 2023), InstructPipe (Zhou et al., 2025), DialogueLab (Hu et al., 2025a), and Sensible Agent (Lee et al., 2025) co-authors, which greatly inspired this project along the way. We appreciate Tim Herrmann and Andrew Helton’s thoughtful reviews. We thank Maryam Sanglaji, Max Spear, Adarsh Kowdle, and Guru Somadder, Shahram Izadi for the directional feedback and contribution.References
@@
- Bez ([n. d.]) [n. d.]. Bezi | AI Assistance for Unity Developers & Studios. https://www.bezi.com/.
@@
- Godot (2022) Godot. 2022. Godot Engine. https://godotengine.org/
@@
- Numan et al. (2023) Nels Numan, Daniele Giunchi, Benjamin Congdon, and Anthony Steed. 2023. Ubiq-Genie: Leveraging External Frameworks for Enhanced Social VR Experiences. In 2023 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, Shanghai, China, 497–501. doi:10.1109/VRW58643.2023.00108